OpenStudy (anonymous):

Pls help urgent help needed! What annual interest rate compounded annually is required to add 50% to an investment in exactly 20 years? Write your answer to 2 decimal places.

terenzreignz (terenzreignz):

Well, let's call that interest rate r... And your initial investment a.... So... \[\huge a(1+r)^{20}=1.5a\] Now solve for r.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@terenzreignz so what will the initial investment be and how did u get 1.5

terenzreignz (terenzreignz):

a is the initial investment, right? And at the end of 20 years, it has to increase by 50% so... it has to be \[\huge a + \color{blue}{<50\% \ of \ a >}=a+0.5a=1.5a\]

terenzreignz (terenzreignz):

You can take the first step and do away with the a, by dividing both sides by a... \[\huge (1+r)^n=1.5\]

terenzreignz (terenzreignz):

Sorry, replace n with 20, my bad :)

OpenStudy (anonymous):

thanks again @terenzreignz

terenzreignz (terenzreignz):

No problem :)

OpenStudy (perl):